### v1

# MaIR: A Locality- and Continuity-Preserving Mamba for Image Restoration

Abstract
Recent advancements in Mamba have shown promising results in image restoration. These methods typically flatten 2D images into multiple distinct 1D sequences along rows and columns, process each sequence independently using selective scan operation, and recombine them to form the outputs. However, such a paradigm overlooks two vital aspects: i) the local relationships and spatial continuity inherent in natural images, and ii) the discrepancies among sequences unfolded through totally different ways. To overcome the drawbacks, we explore two problems in Mamba-based restoration methods: i) how to design a scanning strategy preserving both locality and continuity while facilitating restoration, and ii) how to aggregate the distinct sequences unfolded in totally different ways. To address these problems, we propose a novel Mamba-based Image Restoration model (MaIR), which consists of Nested S-shaped Scanning strategy (NSS) and Sequence Shuffle Attention block (SSA). Specifically, NSS preserves locality and continuity of the input images through the stripe-based scanning region and the S-shaped scanning path, respectively. SSA aggregates sequences through calculating attention weights within the corresponding channels of different sequences. Thanks to NSS and SSA, MaIR surpasses 40 baselines across 14 challenging datasets, achieving state-of-the-art performance on the tasks of image super-resolution, denoising, deblurring and dehazing. Our codes will be available after acceptance.
Figure 1: The scanning strategies in existing Mamba-based methods and our proposed method. (a) Vmamba/Vim uses Z-shaped scan path to flatten 2D image into 1D sequences, in which both the locality and continuity of 2D image are disrupted. (b) Zigma utilizes S-shaped path to maintain spatial continuity, while ignores the locality. (c) LocalMamba leverages window-based scanning region to preserve locality. However, the Z-shaped scanning path within and across the windows disrupts the spatial continuity. In contrast, (d) MaIR divides images into multiple non-overlapping stripes, and adopts S-shaped scanning path within and across the stripes, thus simultaneously preserves both locality and continuity.
1 Introduction
Image restoration aims to recover visually appealing high-quality images from given degraded correspondences, e.g., noisy, blurry, and hazy images. In recent years, the methods based on Convolutional Neural Networks (CNNs) and Transformers have significantly advanced image restoration by effectively capturing locality (i.e., fine-grained patterns and correlations in small regions) and continuity (i.e., smooth, gradual transitions across larger areas) inherent in 2D natural images. To be specific, CNNs capture locality and continuity through the elaborately designed small kernels and sliding strides, respectively. Transformers capture them through local window partitions and adjacent window communications (e.g., window shifts and window expansions). However, like a coin with two sides, the success of CNNs and Transformers in preserving locality and continuity comes at the cost of their ability to capture long-range dependencies. Both of them only consider a limited region of the input image at a time due to their localized kernels or windows, making them challenging to model relationships that span across larger sections of the image. Therefore, it is highly expected to develop a method that is able to capture long-range dependencies while well preserving locality and continuity inherent in 2D natural images.
Mamba [16, 12], a novel selective State Space Model [17], has garnered significant attention due to its promising performance in long sequence modeling while maintaining nearly linear complexity. As Mamba’s core algorithm, Selective Scan Operation (SSO), is inherently designed for 1D sequences, it can not be directly applicable to processing 2D images. To address the problem, Mamba-based restoration methods typically involve a 3-step pipeline: i) flattening 2D image into multiple 1D sequences along rows and columns; ii) processing each sequence independently using SSO; and iii) aggregating the processed sequences to form the output 2D image. However, such a paradigm still faces two demerits when processing images. First, when transforming image into sequences, it disrupts the locality and continuity inherent in image, as illustrated in Fig. 1(a)-(c). Second, it generally aggregates processed sequences via pixel-wise summation, overlooking the distinct contexts among sequences unfolded through totally different ways.
In this work, we present a novel locality- and continuity-preserving Mamba for Image Restoration (MaIR), which consists of Nested S-shaped Scanning strategy (NSS) and Sequence Shuffle Attention block (SSA). Specifically, NSS preserves the locality through stripe-based scanning region, and the continuity via the S-shaped scanning path with shift-stripe mechanism. SSA aggregates the processed sequences by calculating attention weights within corresponding channels of sequences. Thanks to corporation of NSS and SSA, MaIR enjoys the following merits. Firstly, MaIR involves a cost-free solution to preserve the locality and continuity inherent in natural images, ensuring structural coherence and avoiding computational overhead. Secondly, MaIR captures complex dependencies across distinct sequences, facilitating to leverage complementary information from both forward and reversed rows and columns.
To summarize, the contributions and innovations of this work are as below:
  • In this work, we present MaIR, an approach that efficiently captures long-range dependencies while preserving the locality and continuity inherent in natural images.
  • For Mamba, we introduce NSS, a cost-free solution to preserve locality and continuity, and SSA, a module to capture dependencies across distinct sequences.
  • MaIR obtains state-of-the-art performance on four tasks across 14 benchmarks comparing with 40 baselines.
2 Related Works
In this section, we will briefly review related works in image restoration and vision Mamba.
2.1 Image Restoration
According to the focus of this paper, existing methods can be classified into three categories, i.e., CNN-, Transformer- and Mamba-based methods. We will introduce the first two categories here, while the last one is detailed in Sec. 2.2.
CNN-based Method: Benefiting from the ability of capturing locality and continuity in natural images, CNN-based methods have achieved promising results in various tasks of image restoration, such as image super-resolution [29, 64, 11, 40, 27], image denoising [57, 44, 65, 15, 26] and image deblurring [39, 39, 48, 60]. However, since their localized receptive fields, CNNs are inherently limited in capturing long-range dependencies.
Transformer-based Method: Transformers are theoretically capable of capturing the global dependencies [66, 53]. However, to avoid impractical quadratic complexity on images, existing methods [33, 28, 7] tend to partition the local regions of input image into different windows, and calculate attentions within or across the windows. For instance, SwinIR [28] computes attentions within local windows and shifts these windows between layers. HAT [7] divides images into overlapping windows to enhance the interaction between neighbor windows. Although these methods have ensured structural coherence (i.e., locality and continuity) of natural images and avoided computational overhead, they fell into another dilemma of failing to fully capture long-range dependencies due to their limited window sizes.
2.2 Vision Mamba
Due to Mamba’s demonstrated superiority in long-sequence modeling [17, 46, 9], some studies have introduced it into high- [32, 70, 22] and low-level [19, 13, 67] vision tasks. To enable SSO to process images, these methods [32, 70] tend to flatten 2D images into multiple 1D sequences along the different directions. For instance, Vmamba [32] proposes cross-scan strategy which flattens input images along rows and columns. However, existing scanning strategies disrupt structure coherence which is essential for image restoration. Recently, some Mamba-based restoration methods have begun to recognize the importance of structure coherence, and tend to introduce extra coherence-preserving modules. For instance, MambaIR [19] and UVM-Net [67] enhances locality through additional CNN layers, but introduces extra computational costs. Although some other studies [22, 20] devote to designing scanning strategy to preserve locality and continuity, most of them can only preserve one of them. In contrast, MaIR provides a cost-free solution to preserve both locality and continuity.
Figure 2: Illustrations of MaIR. (a) The overall architecture of MaIR, highlighting its core component, Residual Mamba Group (RMG). RMG is primarily composed of (b) Residual Mamba Block (RMB), in which (c) Visual Mamba Module (VMM) plays a pivotal role.
Figure 3: Illustrations of (a) Nested S-shaped Scanning strategy (NSS) and (b) shift-stripe mechanism.
3 Methods
In this section, we first introduce the overall architecture of our MaIR, and then elaborate on NSS and SSA assembled in MaIR Module (MaIRM).
3.1 Overall Architecture
Network Structure: Following previous works [28, 19], MaIR is built up with three stages, namely, shallow feature extraction stage, deep feature extraction stage and reconstruction stage. Specifically, in the shallow feature extraction stage, for a given degraded image $x\in\mathcal{R}^{3\times H\times W}$, we first employ a convolution layer to extract shallow feature $F_{S}\in\mathcal{R}^{C\times H\times W}$, where $H$ and $W$ represent the height and width of $x$, and $C$ is the number of channels. After that, $F_{S}$ is fed to the deep feature extraction stage to produce deep feature $F_{D}\in\mathcal{R}^{C\times H\times W}$. As illustrated in Fig. 2, the deep feature extraction stage is stacked by multiple Residual Mamba Groups (RMGs), where each RMG consists of several Residual Mamba Blocks (RMBs). Within each RMB, a Visual Mamba Module (VMM) is introduced to capture long-range dependencies, which is further composed of our proposed MaIRM. Finally, we reconstruct the high-quality image based on $F_{S}$ and $F_{D}$. Specifically, for image super-resolution, we introduce a pixel-shuffle layer $U_{ps}(\cdot)$ and a $3\times 3$ convolution layer $\Phi_{3\times 3}(\cdot)$ to reconstruct the high-resolution image $y^{\prime}=\Phi_{3\times 3}(U_{ps}(F_{S}+F_{D}))$. For tasks that do not require upsampling (e.g., denoising, deblurring and dehazing), we employ single convolution layer with residual connection to construct high-quality result, which can be formulated as $y^{\prime}=\Phi_{3\times 3}(F_{S}+F_{D})+x$.
Loss Function: For image super-resolution, we use $L_{1}$ loss to optimize the network following [69, 28, 19], which can be formulated as
$\mathcal{L}=\|y-y^{\prime}\|_{1},$
where $y$ is the target image. For image denoising, deblurring and dehazing, we adopt Charbonnier loss, i.e.,
$\mathcal{L}=\sqrt{\|y-y^{\prime}\|^{2}+\epsilon^{2}},$
where $\epsilon$ is a hyper-parameter and set to $10^{-3}$ empirically.
3.2 MaIR Module
As elaborated above, MaIRM serves as the core module of MaIR, which involves a three-step pipeline. To be specific, MaIRM first flattens 2D features into four 1D sequences through NSS along four distinct directions following [32]. Then, MaIRM employs SSO to capture long-range dependencies. Finally, MaIRM aggregates processed sequences through SSA to form outputs. Mathematically, for input feature $F_{i,j}$, output feature $F^{M}_{i,j}$ can be formulated as
$F^{M}_{i,j}$ $=M_{i,j}(F_{i,j}),$
$=\Phi_{i,j}^{SSA}(\Phi^{SSO}_{i,j}(\Phi_{i,j}^{NSS}(F_{i,j}))),$
where $M_{i,j}(\cdot)$, $\Phi_{i,j}^{NSS}(\cdot)$, $\Phi_{i,j}^{SSO}(\cdot)$ and $\Phi_{i,j}^{SSA}(\cdot)$ are MaIRM, NSS, SSO and SSA in the $j$-th RMB of the $i$-th RMG, respectively.
NSS: NSS is designed to extract locality- and continuity-preserving sequences from input features. Motivated by the observation illustrated in Fig. 1, one could find that i) LocalMamba [22] preserves locality through restricted scanning region, and ii) Zigma [20] preserves continuity through S-shaped scanning path. Thus, as shown in Fig. 3(a), we design the nested S-shaped scanning strategy, which divides features into multiple non-overlapping stripes and uses S-shaped scanning path within and across stripes to maintain both locality and continuity. To better leverage spatial information, we extract sequences with four different scanning directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, following previous works [32, 19].
Besides, NSS includes shift-stipe mechanism to preserve locality and continuity on the boundary regions between adjacent stripes. As depicted in Fig. 3(b), for two successive modules, the first module partitions features into multiple non-overlapping stripes with stripe width $w_{s}$. For the second module, we employ the shift-stripe operation, and set the first and last stripe widths as $\frac{w_{s}}{2}$ and others’ width as $w_{s}$. Consequently, the boundary regions in the previous module will be fully covered by a single stripe in this module.
Figure 4: Illustration of the Sequence Shuffle Attention (SSA). The input features $\{X^{i}\}_{i=1}^{K}\in\mathcal{R}^{D\times H\times W}$ are first pooled and concatenated to form $\tilde{X}\in\mathcal{R}^{L}$, where $L=K\times D$. This sequence undergoes the sequence shuffle operation and results in shuffled sequences $\hat{X}\in\mathcal{R}^{L}$, whose channels are split by $D$ group. Then, group convolution and sequence unshuffle operation are applied, producing unshuffled weights $\tilde{W}\in\mathcal{R}^{L}$, which are further chunked and reshaped into attention weights $\{W^{i}\}_{i=1}^{K}\in\mathcal{R}^{D}$. Finally, the output feature $Y\in\mathcal{R}^{D\times H\times W}$ is computed by performing a weighted summation of the input features using the attention weights.
SSA: SSA aggregates the processed sequences by calculating attentions within corresponding channels. This design enables it to capture complex dependencies across distinct sequences, thus better leveraging complementary information from different scanning directions.
As shown in Fig. 4, supposing sequence number $K=4$, for SSO-processed sequences $\{X^{i}\}_{i=1}^{4}$, we first apply spatial average pooling $\Phi_{AP}(\cdot)$ to reduce the computational cost, and then concatenate as
$\tilde{X}$ $=\Phi_{cat}(\Phi_{AP}(\{X^{i}\}_{i=1}^{4}))$
$=[x^{1}_{1},\cdots,x^{1}_{D},x^{2}_{1},\cdots,x^{2}_{D},x^{3}_{1},\cdots,x^{3}_{D},x^{4}_{1},\cdots,x^{4}_{D}],$
where $x^{k}_{d}$ is the pooled feature in $d$-th channel of $k$-th sequence, and $D$ is the number of channel in MaIRM. Then, we employ sequence shuffle operation $\Phi_{ss}(\cdot)$ to rearrange features into
$\hat{X}$ $=\Phi_{ss}(X)$
$=[x^{1}_{1},x^{2}_{1},x^{3}_{1},x^{4}_{1},x^{1}_{2},x^{2}_{2},x^{3}_{2},x^{4}_{2},\cdots,x^{1}_{D},x^{2}_{D},x^{3}_{D},x^{4}_{D}].$
After that, we employ group convolution $\Phi_{g}(\cdot)$ with group size four to obtain the channel-wise attention weights and unshuffle the weights back to their original order, i.e.,
$\tilde{W}$ $=\Phi_{su}(\Phi_{g}(\hat{X}))$
$=[w^{1}_{1},\cdots,w^{1}_{D},w^{2}_{1},\cdots,w^{2}_{D},w^{3}_{1},\cdots,w^{3}_{D},w^{4}_{1},\cdots,w^{4}_{D}],$
where $\Phi_{su}(\cdot)$ is sequence unshuffle operation. The unshuffled weights $\tilde{W}$ are chunked as $\{W^{i}\}_{i=1}^{4}=\Phi_{chunk}(\tilde{W}),$ where $\Phi_{chunk}(\cdot)$ refers to the chunk operation. Finally, we adopt weight summation based on $\{W^{i}\}_{i=1}^{4}$ to generate the output, which can be formulated as:
$Y$ $=\sum_{i=1}^{K=4}W^{i}*X^{i},$
and $Y$ is the output sequence of SSA.
4 Experiments
In this section, we evaluate our MaIR on four representative image restoration tasks, i.e., image super-resolution, image denoising, image deblurring, and image dehazing. In the following, we will first show quantitative results, and then conduct analysis studies to verify the reasonability. Experimental settings will be presented in the supplementary materials.
Table 1: Quantitative results on classic image super-resolution. The best and second best results are in red and blue.
Methods Scale Set5 Set14 B100 Urban100 Manga109
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SAN [11] ×$\times 2$ 38.31 0.9620 34.07 0.9213 32.42 0.9028 33.10 0.9370 39.32 0.9792
HAN [40] ×$\times 2$ 38.27 0.9614 34.16 0.9217 32.41 0.9027 33.35 0.9385 39.46 0.9785
IGNN [68] ×$\times 2$ 38.24 0.9613 34.07 0.9217 32.41 0.9025 33.23 0.9383 39.35 0.9786
NLSA [38] ×$\times 2$ 38.34 0.9618 34.08 0.9231 32.43 0.9027 33.42 0.9394 39.59 0.9789
ELAN [63] ×$\times 2$ 38.36 0.9620 34.20 0.9228 32.45 0.9030 33.44 0.9391 39.62 0.9793
IPT [4] ×$\times 2$ 38.37 - 34.43 - 32.48 - 33.76 - - -
SwinIR [28] ×$\times 2$ 38.42 0.9623 34.46 0.9250 32.53 0.9041 33.81 0.9427 39.92 0.9797
SRFormer [69] ×$\times 2$ 38.51 0.9627 34.44 0.9253 32.57 0.9046 34.09 0.9449 40.07 0.9802
MambaIR [19] ×$\times 2$ 38.57 0.9627 34.67 0.9261 32.58 0.9048 34.15 0.9446 40.28 0.9806
MaIR ×$\times 2$ 38.56 0.9628 34.75 0.9268 32.59 0.9049 34.19 0.9451 40.30 0.9807
MaIR+ ×$\times 2$ 38.62 0.9630 34.82 0.9272 32.62 0.9053 34.38 0.9462 40.48 0.9811
SAN [11] ×$\times 3$ 34.75 0.9300 30.59 0.8476 29.33 0.8112 28.93 0.8671 34.30 0.9494
HAN [40] ×$\times 3$ 34.75 0.9299 30.67 0.8483 29.32 0.8110 29.10 0.8705 34.48 0.9500
IGNN [68] ×$\times 3$ 34.72 0.9298 30.66 0.8484 29.31 0.8105 29.03 0.8696 34.39 0.9496
NLSA [38] ×$\times 3$ 34.85 0.9306 30.70 0.8485 29.34 0.8117 29.25 0.8726 34.57 0.9508
ELAN [63] ×$\times 3$ 34.90 0.9313 30.80 0.8504 29.38 0.8124 29.32 0.8745 34.73 0.9517
IPT [4] ×$\times 3$ 34.81 - 30.85 - 29.38 - 29.49 - - -
SwinIR [28] ×$\times 3$ 34.97 0.9318 30.93 0.8534 29.46 0.8145 29.45 0.8826 35.12 0.9537
SRFormer [69] ×$\times 3$ 35.02 0.9323 30.94 0.8540 29.48 0.8156 30.04 0.8865 35.26 0.9543
MambaIR [19] ×$\times 3$ 35.08 0.9323 30.99 0.8536 29.51 0.8157 29.93 0.8841 35.43 0.9546
MaIR ×$\times 3$ 35.10 0.9324 31.05 0.8541 29.51 0.8160 30.05 0.8863 35.44 0.9547
MaIR+ ×$\times 3$ 35.15 0.9328 31.12 0.8550 29.56 0.8167 30.24 0.8881 35.67 0.9556
SAN [11] ×$\times 4$ 32.64 0.9003 28.92 0.7888 27.78 0.7436 26.79 0.8068 31.18 0.9169
HAN [40] ×$\times 4$ 32.64 0.9002 28.90 0.7890 27.80 0.7442 26.85 0.8094 31.42 0.9177
IGNN [68] ×$\times 4$ 32.57 0.8998 28.85 0.7891 27.77 0.7434 26.84 0.8090 31.28 0.9182
NLSA [38] ×$\times 4$ 32.59 0.9000 28.87 0.7891 27.78 0.7444 26.96 0.8109 31.27 0.9184
ELAN [63] ×$\times 4$ 32.75 0.9022 28.96 0.7914 27.83 0.7459 27.13 0.8167 31.68 0.9226
IPT [4] ×$\times 4$ 32.64 - 29.01 - 27.82 - 27.26 - - -
SwinIR [28] ×$\times 4$ 32.92 0.9044 29.09 0.7950 27.92 0.7489 27.45 0.8254 32.03 0.9260
SRFormer [69] ×$\times 4$ 32.93 0.9041 29.08 0.7953 27.94 0.7502 27.68 0.8311 32.21 0.9271
MambaIR [19] ×$\times 4$ 33.03 0.9046 29.20 0.7961 27.98 0.7503 27.68 0.8287 32.32 0.9272
MaIR ×$\times 4$ 32.93 0.9045 29.20 0.7958 27.98 0.7507 27.71 0.8305 32.46 0.9284
MaIR+ ×$\times 4$ 33.14 0.9058 29.28 0.7974 28.02 0.7516 27.89 0.8336 32.66 0.9297
Table 2: Quantitative results on lightweight image super-resolution. The best and second best results are in red and blue.
Methods Scale Params MACs Set5 Set14 B100 Urban100 Manga109
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
CARN [2] ×$\times 2$ 1,592K 222.8G 37.76 0.9590 33.52 0.9166 32.09 0.8978 31.92 0.9256 38.36 0.9765
IMDN [23] ×$\times 2$ 694K 158.8G 38.00 0.9605 33.63 0.9177 32.19 0.8996 32.17 0.9283 38.88 0.9774
LAPAR-A [27] ×$\times 2$ 548K 171.0G 38.01 0.9605 33.62 0.9183 32.19 0.8999 32.10 0.9283 38.67 0.9772
LatticeNet [34] ×$\times 2$ 756K 169.5G 38.15 0.9610 33.78 0.9193 32.25 0.9005 32.43 0.9302 - -
SwinIR [28] ×$\times 2$ 910K 122.2G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
MambaIR-Tiny [19] ×$\times 2$ 905K 167.1G 38.13 0.9610 33.95 0.9208 32.31 0.9013 32.85 0.9349 39.20 0.9782
MaIR-Tiny ×$\times 2$ 878K 207.8G 38.18 0.9610 33.89 0.9209 32.31 0.9013 32.89 0.9346 39.22 0.9778
MambaIR-Small [19] ×$\times 2$ 1,363K 567.5G 38.16 0.9610 34.00 0.9212 32.34 0.9017 32.92 0.9356 39.31 0.9779
MaIR-Small ×$\times 2$ 1,355K 542.0G 38.20 0.9611 33.91 0.9209 32.34 0.9016 32.97 0.9359 39.32 0.9779
CARN [2] ×$\times 3$ 1,592K 111.8G 34.29 0.9255 30.29 0.8407 29.06 0.8034 28.06 0.8493 33.50 0.9440
IMDN [23] ×$\times 3$ 703K 71.5G 34.36 0.9270 30.32 0.8417 29.09 0.8046 28.17 0.8519 33.61 0.9445
LAPAR-A [27] ×$\times 3$ 544K 144.0G 34.36 0.9267 30.34 0.8421 29.11 0.8054 28.15 0.8523 33.51 0.9441
LatticeNet [34] ×$\times 3$ 765K 76.3G 34.53 0.9281 30.39 0.8424 29.15 0.8059 28.33 0.8538 - -
SwinIR [28] ×$\times 3$ 918K 55.4G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
MambaIR-Tiny [19] ×$\times 3$ 913K 74.5G 34.63 0.9288 30.54 0.8459 29.23 0.8084 28.70 0.8631 34.12 0.9479
MaIR-Tiny ×$\times 3$ 886K 93.0G 34.68 0.9292 30.54 0.8461 29.25 0.8088 28.83 0.8651 34.21 0.9484
MambaIR-Small [19] ×$\times 3$ 1,371K 252.7G 34.72 0.9296 30.63 0.8475 29.29 0.8099 29.00 0.8689 34.39 0.9495
MaIR-Small ×$\times 3$ 1,363K 241.4G 34.75 0.9300 30.63 0.8479 29.29 0.8103 28.92 0.8676 34.46 0.9497
CARN [2] ×$\times 4$ 1,592K 90.9G 32.13 0.8937 28.60 0.7806 27.58 0.7349 26.07 0.7837 30.47 0.9084
IMDN [23] ×$\times 4$ 715K 40.9G 32.21 0.8948 28.58 0.7811 27.56 0.7353 26.04 0.7838 30.45 0.9075
LAPAR-A [27] ×$\times 4$ 659K 94.0G 32.15 0.8944 28.61 0.7818 27.61 0.7366 26.14 0.7871 30.42 0.9074
LatticeNet [34] ×$\times 4$  [^7] K 43.6G 32.30 0.8962 28.68 0.7830 27.62 0.7367 26.25 0.7873 - -
SwinIR [28] ×$\times 4$ 930K 31.8G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.48 0.7980 30.92 0.9151
MambaIR-Tiny [19] ×$\times 4$ 924K 42.3G 32.42 0.8977 28.74 0.7847 27.68 0.7400 26.52 0.7983 30.94 0.9135
MaIR-Tiny ×$\times 4$ 897K 53.1G 32.48 0.8985 28.81 0.7864 27.71 0.7414 26.60 0.8013 31.13 0.9161
MambaIR-Small [19] ×$\times 4$ 1,383K 143.0G 32.51 0.8993 28.85 0.7876 27.75 0.7423 26.75 0.8051 31.26 0.9175
MaIR-Small ×$\times 4$ 1,374K 136.6G 32.62 0.8998 28.90 0.7882 27.77 0.7431 26.73 0.8049 31.34 0.9183
Figure 5: Visual comparison of ×$\times 4$ image super-resolution results on the Manga109 dataset. MaIR demonstrates superior visual quality, particularly in preserving fine details and textures.
4.1 Results on Image Super-Resolution
In this section, we conduct experiments on both classic and lightweight image super-resolution.
Datasets: Following previous works [28, 19], we employ DF2K (DIV2K [49]+Flickr2K [30]) as the training set for classic image super-resolution, and DIV2K as training set for lightweight image super-resolution. For evaluation, we employ the following five datasets as test sets, i.e., Set5 [3], Set14 [54], B100 [36], Urban100 [21] and Manga109 [37]. Following existing works [64, 58, 11, 40], the low-resolution images are downsampled from the corresponding high-resolution images via bicubic interpolation.
Baselines: We compare our method with 15 competitive baselines. Specifically, we adopt four CNN-based methods (i.e., SAN [11], HAN [40], IGNN [68], and NLSA [38]), four transformer-based methods (i.e., ELAN [63], IPT [4], SwinIR [28] and SRFormer [69]) and one Mamba-based method (i.e., MambaIR [19]) as the baselines for classic super-resolution. For lightweight super-resolution, four CNN-based methods (i.e., CARN [2], IMDN [23], LAPAR [27], LatticeNet [34]), two transformer-based methods (i.e., SwinIR [28] and SRFormer [69]) and one Mamba-based method (i.e., MambaIR [19]) are introduced into both quantitative and qualitative comparisons. Similar to MambaIR, which offers two versions for lightweight super-resolution, MaIR is also available in two configurations: MaIR-Tiny and MaIR-Small.
Results: For classic super-resolution, as shown in Tabs. 1 and 5, one could observe that MaIR achieves the best result in almost all quantitative comparisons. For instance, our method surpasses MambaIR [19] with 0.03dBsimilar-$\sim$0.12dB in terms of PSNR on Urban100, and SRFormer with at most 0.04dB, 0.10dB and 0.25dB in terms of PSNR on B100, Urban100, and Manga109, respectively, which demonstrates the superiority of MaIR. For light-weight SR, MaIR also exhibits its advancement compared to baselines as reported in Tab. 2. Taking ×$\times 4$ scale as examples, MaIR-Small surpasses MambaIR-Small by 0.08dB in terms of PSNR on Manga109 with fewer parameters and MACs. MaIR-Tiny outperforms MambaIR-Tiny and SwinIR by 0.08dB and 0.12dB in terms of PSNR on Urban100 with fewer parameters, which verifies both the efficiency and effectiveness of our proposed method.
Table 3: Quantitative results on gaussian color image denoising. The best and second best results are in red and blue.
Methods BSD68 Kodak24 McMaster Urban100
$\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50
IRCNN [58] 33.86 31.16 27.86 34.69 32.18 28.93 34.58 32.18 28.91 33.78 31.20 27.70
FFDNet [59] 33.87 31.21 27.96 34.63 32.13 28.98 34.66 32.35 29.18 33.83 31.40 28.05
DnCNN [57] 33.90 31.24 27.95 34.60 32.14 28.95 33.45 31.52 28.62 32.98 30.81 27.59
DRUNet [61] 34.30 31.69 28.51 35.31 32.89 29.86 35.40 33.14 30.08 34.81 32.60 29.61
SwinIR [28] 34.42 31.78 28.56 35.34 32.89 29.79 35.61 33.20 30.22 35.13 32.90 29.82
Restormer [53] 34.40 31.79 28.60 35.47 33.04 30.01 35.61 33.34 30.30 35.13 32.96 30.02
CODE [66] 34.33 31.69 28.47 35.32 32.88 29.82 35.38 33.11 30.03 - - -
ART [56] 34.46 31.84 28.63 35.39 32.95 29.87 35.68 33.41 30.31 35.29 33.14 30.19
MambaIR [19] 34.43 31.80 28.61 35.34 32.91 29.85 35.62 33.35 30.31 35.17 32.99 30.07
MaIR 34.48 31.86 28.66 35.53 33.09 30.04 35.71 33.44 30.35 35.35 33.22 30.30
MaIR+ 34.50 31.88 28.69 35.56 33.13 30.08 35.74 33.48 30.39 35.42 33.30 30.41
Table 4: Quantitative results on real image denoising. The best and second best results are in red and blue.
DeamNet [43] MPRNet [52] NBNet [8] DAGL [10] Uformer [50] MambaIR [19] MaIR
PSNR 39.47 39.71 39.75 38.94 39.89 39.89 39.92
SSIM 0.957 0.958 0.959 0.953 0.960 0.960 0.960
Figure 6: Visual comparison of image denoising results on the Urban100 dataset. MaIR effectively removes noise in the images and produces detailed textures that closely match the ground truth.
4.2 Results on Image Denoising
In this section, we evaluate MaIR on both synthetic Gaussian noise and real-world noise.
Datasets: For synthetic noise removal, we train MaIR on DFWB, which consists of DIV2K, Flickr2K, Waterloo Exploration Dataset (WED) [35] and BSD400 [36]. For evaluation, we utilize BSD68 [36], Kodak24, McMaster [62], and Urban100 as test set. Following [57, 59, 58, 28], we generate noisy images by manually adding white Gaussian noise to the clean images with three distinct noise levels, i.e., $\sigma=15,25,50$. For real-world image denoising, our model is trained and tested on the SIDD-Medium [1] dataset, which provides 320 high-resolution noisy-clean image pairs for training and additional 40 image pairs for test.
Baselines: We compare our MaIR with 14 representative methods. To be specific, we adopt four CNN-based methods (i.e., IRCNN [58], FFDNet [59], DnCNN [57] and DRUNet [61]), four transformer-based methods (i.e., SwinIR [28], Restormer [53], CODE [66] and ART [56]) and one Mamba-based method (i.e., MambaIR [19]) as the baselines for synthetic noise removal. For real-world image denoising, four CNN-based methods (i.e., DeamNet [43], MPRNet [52], NBNet [8] and DAGL [10]), two transformer-based methods (i.e., Uformer [50] and Restormer [53]) and one Mamba-based method (i.e., MambaIR [19]) are introduced for comparisons.
Results: As depicted in the Tab. 3-4, MaIR demonstrates superior performance on both synthetic and real-world image denoising compared to baselines. Taking results on Urban100 as examples, MaIR averagely outperforms MambaIR by 0.21dB in terms of PSNR, indicates its superiority on image denoising. Similar results can be derived from the qualitative comparisons shown in Fig. 6, MaIR could keep more detailed textures on the restored images, which are more closely to the ground truth.
Table 5: Quantitative results on image motion deblurring. The best and second best results are in red and blue. MACs in this table are evaluated on 128$\times$128 patches followed [66].
Method Params MACs GoPro HIDE
SRN [48] 3.76M 35.87G 30.26 28.36
DBGAN [60] 11.59M 379.92G 31.10 28.94
MT-RNN [41] 2.64M 13.72G 31.15 29.15
DMPHN [55] 86.80M - 31.20 29.09
CODE [66] 12.18M 22.52G 31.94 29.67
MIMO+ [9] 16.10M 38.64G 32.45 29.99
MPRNet [52] 20.13M 194.42G 32.66 30.96
Restormer [53] 26.13M 35.31G 32.92 31.22
Uformer [50] 50.88M 22.36G 33.06 30.90
CU-Mamba [13] 19.7M - 33.53 31.47
NAFNet [6] 67.89M 15.85G 33.69 31.32
MaIR 26.29M 49.29G 33.69 31.57
Figure 7: Visual comparison of motion deblurring results on the GoPro dataset. MaIR demonstrates superior performance in effectively removing motion blur while preserving precise fine details and textures, closely matching the ground truth.
4.3 Results on Image Deblurring
In this section, we evaluate MaIR on motion deblurring to verify the effectiveness of our proposed method.
Datasets: Following previous works [52, 53], we employ GoPro dataset [39] for training which consists of 2,103 blurry-clean image pairs. For evaluation, we use two common datasets, i.e., GoPro test set and HIDE [45], which consist of 1,111 and 2,025 blurry-clean pairs, respectively.
Baselines: We adopt 11 competitive image deblurring baselines for comparisons. In detail, we adopt six CNN-based deblurring methods (i.e., SRN [48], DBGAN [60], DMPHN [55], MIMO [9], MPRNet [52], and NAFNet [6]), three transformer-based methods (i.e., CODE [66], Restormer [53] and Uformer [50]), one RNN-based method (i.e., MT-RNN [41]) and one Mamba-based method (i.e., CU-Mamba [13]) as the baselines.
Results: As shown in Tab. 5, proposed MaIR surpasses other baselines by PSNR on both GoPro and HIDE. In detail, MaIR outperforms Restormer [53] by 0.77dB on the GoPro dataset and by 0.35dB on the HIDE dataset in terms of PSNR. Although NAFNet achieves similar quantitative results on GoPro, MaIR surpasses NAFNet on HIDE dataset by 0.25dB in terms of PSNR. As illustrated in Fig. 7, MaIR demonstrates its ability on handling heavily degraded areas, i.e., MaIR could effectively remove blur and restore details around the wheels.
Table 6: Quantitative results on image dehazing. The best and second best results are in red and blue. MACs in this table are evaluated on 256$\times$256 patches followed [47, 67].
Method AODNet GDN MSBDN FFANet AECRNet Dehamer DehazeFormer UVM-Net MaIR
[24] [31] [14] [42] [51] [18] [47] [67] (Ours)
Params 0.002M 0.96M 31.35M 4.46M 2.61M 132.45M 4.63M 1,003.94M 3.40M
MACs 0.115G 21.49G 41.54G 287.8G 52.20G 48.93G 48.64G 501.91G 24.03G
SOTS- PSNR 20.51 32.16 33.67 36.39 37.17 36.63 38.46 40.17 39.45
Indoor SSIM 0.816 0.984 0.985 0.989 0.990 0.988 0.994 0.996 0.997
SOTS- PSNR 24.14 30.86 33.48 33.57 - 35.18 34.29 34.92 36.96
Outdoor SSIM 0.920 0.982 0.982 0.984 - 0.986 0.983 0.984 0.991
SOTS- PSNR 20.27 25.86 28.56 29.96 28.52 - 30.89 31.92 31.52
Mix SSIM 0.855 0.944 0.966 0.973 0.964 - 0.977 0.982 0.980
Figure 8: Visual comparison of image dehazing results on the SOTS dataset. MaIR can effectively remove haze and restore content with colors that closely match the ground truth.
4.4 Results on Image Dehazing
In this section, we evaluate MaIR on image dehazing to verify the effectiveness of MaIR.
Datasets: Following existing works [47], we employ RESIDE dataset [25] for training and testing. For indoor scenes, we train MaIR on Indoor Training Set (ITS) which consists of 13,990 hazy-clean pairs, and test it on indoor synthetic objective testing set (SOTS-Indoor) involving 500 pairs. For outdoor scenes, we train MaIR on Outdoor Training Set (OTS), which contains 313,950 image pairs, and evaluate it on outdoor synthetic objective testing set (SOTS-Outdoor) involving 500 images. In addition, to verify MaIR on more general cases, we also train the model on RESIDE-6K and test it on the SOTS-mix, which mix both indoor and outdoor images.
Baselines: We adopt eight competitive methods as baselines. Specifically, we adopt five CNN-based image dehazing methods (i.e., AODNet [24], GDN [31], MSBDN [14], FFANet [42] and AECRNet [51]), two transformer-based methods (i.e., Dehamer [18], Dehazeformer [47]) and one Mamba-based method (i.e., UVM-Net [67]) as baselines.
Results: As shown in Tabs. 6 and 8, our MaIR surpasses most baselines on both quantitative and qualitative comparisons. Taking quantitative results as examples, MaIR significantly outperforms DehazeFormer and UVM-Net by 2.67dB and 2.04dB in terms of PSNR on the outdoor scenes. Although UVM-Net is slightly higher on PSNR in the indoor scenes, MaIR only takes 0.3% and 4.8% params and MACs of the UVM-Net, which verifies both effectiveness and efficiency.
4.5 Analysis Experiments
In this section, we first conduct ablation studies to verify the effectiveness of the NSS and SSA. Then, we introduce analysis experiments to investigate the impact of stripe width on the overall performance.
Table 7: Ablation study on the proposed NSS scheme, tested on lightweight super-resolution tasks with scale factor ×$\times 2$. The results on the Urban100 dataset are presented, which demonstrates the effectiveness of the NSS.
Baseline w/o NSS w/o SS LM ZigMa Hilbert
PSNR 32.97 32.94 32.93 32.93 32.88 32.95
SSIM 0.9359 0.9355 0.9351 0.9357 0.9354 0.9356
4.5.1 Ablation Studies
We first conduct ablation study to analyze the effectiveness of NSS. In detail, five configurations are conducted, i) replacing NSS with Z-shaped scanning strategy (denoted as w/o NSS), ii) removing shift stripe (denoted as w/o SS), iii) replacing NSS with the scanning strategy in LocalMamba [22] (denoted as LM), iv) replacing NSS by the scanning strategy in ZigMa [20] (denoted as ZigMa) and v) replacing NSS by the Hilbert curve (denoted as Hilbert). As illustrated in Tab. 7, one could see that NSS is important to improve the performance of MaIR.
Table 8: Ablation study on SSA for lightweight super-resolution with scale factor ×$\times 2$. The results on Urban100 dataset are presented, which demonstrates the effectiveness of the NSS.
Baseline w.o. SSA Gating
PSNR 32.97 32.90 32.92
SSIM 0.9359 0.9350 0.9252
To investigate the effectiveness of SSA, we also conduct experiments under the following settings: i) removing proposed SSA and aggregating sequences via sequences-wise addition (termed as w.o. SSA), and ii) replacing SSA by the gating strategy in [5] (termed as Gating). As reported in Tab. 8, SSA is more effective than others.
Table 9: Analyses on stripe widths. Experiment was conducted on the Urban100 dataset with a scale factor ×$\times 2$ for lightweight super-resolution tasks, which illustrates how changes in stripe width affect the restored image quality.
2 4 8 16 32
PSNR 32.95 32.97 32.97 32.97 32.92
SSIM 0.9356 0.9359 0.9355 0.9357 0.9355
4.5.2 Results on different stripe width
To investigate influence of stripe width, we train lightweight SR model with various stripe width (2, 4, 8, 16 and 32) and evaluate them on Urban100 dataset. As presented in Tab. 9, one could observe that the PSNR and SSIM values are quite similar under different settings, except for the cases with the largest and the smallest stripe widths. It indicates that the proposed method exhibits robustness against changes in stripe width, maintaining high-quality image restoration across a range of stripe widths.
5 Conclusion
In this paper, we propose MaIR, a novel state space model for image restoration that can preserve both local dependencies and spatial continuity of input images. To this end, we propose two designs: Nested S-shaped Scanning strategy (NSS) and Sequences Shuffle Attention (SSA). NSS is designed to extract locality- and continuity-preserving sequences from images, and SSA adaptively aggregates these sequences. Thanks to their cooperation, MaIR not only addresses the limitations of existing Mamba-based restoration methods but also improves image quality without introducing extra computations. Extensive experiments across four tasks on 14 benchmarks comparing with 40 baselines, validate the superiority of MaIR, demonstrating its robustness and effectiveness in various image restoration tasks.